How many different samples of size n = 3 can be selected from a population with the following sizes?

1.) N = 4

2.) N = 8

3.) N = 20

4.) N = 50



1
Expert's answer
2022-03-28T13:20:52-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }."


"1.) \ \begin{pmatrix}
 4 \\
 3
\end{pmatrix}=\cfrac{4! } {3! \cdot1! }=4."

"2.) \ \begin{pmatrix}
 8\\
 3
\end{pmatrix}=\cfrac{8! } {3! \cdot5! }=\\
=\cfrac{6\cdot7\cdot8}{2\cdot3}=56."

"3.) \ \begin{pmatrix}
 20\\
 3
\end{pmatrix}=\cfrac{20! } {3! \cdot17! }=\\
=\cfrac{18\cdot19\cdot20}{2\cdot3}=1140."

"4.) \ \begin{pmatrix}
 50\\
 3
\end{pmatrix}=\cfrac{50! } {3! \cdot47! }=\\
=\cfrac{48\cdot49\cdot50}{2\cdot3}=19600."